OpenStudy (anonymous):

Find w(-3) using synthetic substitution w(x)=11x^3-6x^2+2

OpenStudy (anonymous):

Sorry it's w(-3)

OpenStudy (jdoe0001):

hmmm do you know how to do synthetic division?

OpenStudy (anonymous):

A little bit

OpenStudy (jdoe0001):

ok... let us drop our 1st number and multiply by -3 \(\bf x=-3\implies x+3=0\implies (x+3)=0\\ \quad \\ 11x^3-6x^2+2\implies 11x^3-6x^2+0x+2\quad \textit{arrange in descending order}\\ \quad \\ \begin{array}{rrrrrrrl} -3&|&11&-6&+0&+2\\ &|&&-33\\ \hline\\ &&11&-39 \end{array} \)

OpenStudy (jdoe0001):

so -6-33 = -39 so now let's multiply -3 * -39 and use the product in the following term \(\bf \begin{array}{rrrrrrrl} -3&|&11&-6&+0&+2\\ &|&&-33&117\\ \hline\\ &&11&-39&117 \end{array}\)

OpenStudy (jdoe0001):

and now -3 * 117 to be added to +2, the following term \(\bf \begin{array}{rrrrrrrl} -3&|&11&-6&+0&+2\\ &|&&-33&117&-351\\ \hline\\ &&11&-39&117&-349 \end{array}\)

OpenStudy (jdoe0001):

so we use those coefficients, in our new expression lessened by 1 degree from the original the original was 3rd degree, the quotient will be 2nd degree then \(\bf \begin{array}{rrrrrrrl} -3&|&11&-6&+0&+2\\ &|&&-33&117&-351\\ \hline\\ &&11&-39&117&-349 \end{array}\implies 11x^2-39x+117-\frac{-349}{x+3}\)

OpenStudy (jdoe0001):

the last term is the remainder, so it stays in "division form", or you can just say it's the remainder

OpenStudy (anonymous):

Is the last supposed to be - -349? So I would write that like 117- (-349/x+3)

OpenStudy (jdoe0001):

well.... the "remainder" stays as a division, so it'd be written as a fraction

OpenStudy (jdoe0001):

ohh.... ahemm... I see what you mean, well - * - = + .... so could be also written as \(\bf 11x^2-39x+117+\frac{349}{x+3}\) but still as fractional

OpenStudy (anonymous):

Oh ok, but I can't type a fractional in word...

OpenStudy (jdoe0001):

check in MS Word under I think is under ..... Format > insert > Equation Word has an equation editor

OpenStudy (anonymous):

I think the teacher will understand though, how do I do direct substitution for my next problem? Do I just plug in the number?

OpenStudy (jdoe0001):

well.... you question was " Find w(-3) using synthetic substitution" w(-3) will give you the REMAINDER you simply were asked to get it by using synthetic division so the answer I gather is -349 you can check by doing w(-3) on the equation and using direct substitution like \(\bf w(x)=11x^3-6x^2+2\implies w(-3)=11(-3)^3-6(-3)^2+2\)

OpenStudy (jdoe0001):

that should yield you as the synthetic division did, -349

OpenStudy (anonymous):

The direct substitution problem is: Find m(5) using direct substitution: m(x)=x^2+x-1

OpenStudy (anonymous):

Do I just plug 5 into the equation?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

So 5^2+5-1= 29. That's the answer?

OpenStudy (jdoe0001):

yes, that means that 29 is the remainder were you to use synthetic diviision

OpenStudy (anonymous):

And for the synthetic substitution problem I should write out the entire answer right? Like the whole 11x^2 to the remainder

OpenStudy (jdoe0001):

I'd say yes

OpenStudy (anonymous):

I agree. Thanks a lot man! If I have more questions, would you mind if I tag you?

OpenStudy (jdoe0001):

sometimes you may just be asked to write the expression without the remainder and then write the remainder on the side

OpenStudy (jdoe0001):

tis ok